# P2LHAP:Wearable sensor-based human activity recognition, segmentation and forecast through Patch-to-Label Seq2Seq Transformer

### v1

@@ III-B2 Encoder
We use the native Transformer encoder [dosovitskiy2020image] to model the features of patch-level sensor sequences, similar to the method used in [41]. The Encoder module utilizes trainable linear projections to map continuous Patch blocks to a D-dimensional Transformer latent space $\mathbf{W}_{p}\in\mathbb{R}^{D\times P}$. It incorporates a learnable additive positional encoding $\mathbf{W}_{pos}\in\mathbb{R}^{D\times N}$ to capture the temporal order of the patches, represented by $\mathbf{x}_{d}^{(i)}=\mathbf{W}_{p}\mathbf{x}_{p}^{(i)}+\mathbf{W}_{pos}$, where $\mathbf{x}_{d}^{(i)}\in\mathbb{R}^{D\times N}$ represents the input fed into the Transformer encoder. For each head $h=1,\cdots,H$ in the multi-head attention, the Patch blocks are transformed into query matrices $\mathbf{Q}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{Q}$, key matrices $\mathbf{K}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{K}$, and value matrices $\mathbf{V}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{V}$, where $\mathbf{W}_{h}^{Q},\mathbf{W}_{h}^{K}\in\mathbb{R}^{D\times d_{k}},\mathbf{W}_{h}^{V}\in\mathbb{R}^{D\times D}$. Subsequently, the attention outputs are normalized to obtain the final attention output $\mathbf{Q}_{h}^{(i)}\in\mathbb{R}^{D\times N}:$
@@ III-B3 Decoder
$\mathbf{z}^{{}^{\prime}(i)}=softmax(\mathbf{Q}^{T}\mathbf{K}/\tau)\mathbf{V}=f_{\theta}(\mathbf{X}^{{}^{\prime}},\mathbf{Z}^{(i)})$ (3)
Here $\tau$ is the temperature parameter, and $f_{\theta}(\cdot)$ represents the attention mechanism network.
@@ III-C Activity classification
Among them, $\hat{y}$ and $y$ represent the patch-level prediction label and the real label, respectively. Given the imbalance of human activity classes under unconstrained conditions, we employ the Softmax CrossEntropy loss function $l_{p}(\cdot)$ and consider the number of effective samples for classification. The calculation formula is as follows:
@@
Where $L$ signifies the length of input sequence, and C represents the activity class number of dataset. In sensor-based HAR tasks, we leverage this loss function and consider the number of effective samples to amplify the weight assigned to minority classes. This approach aims to yield fairer and more robust results for each individual class.
@@ III-D Activity segmentation and smoothing strategy
Given that the classification loss treats each patch independently, it may result in unwarranted over-segmentation effects. To achieve smoother transitions between patches and minimize over-segmentation errors, we employ the truncated mean square error as our chosen smoothing loss function. The formulation of the smoothing loss is as follows:
@@
For the activity label sequence $\hat{y}$ output by the model, in order to reduce the impact of over-segmentation on subsequent activity merging, we set each activity label as the center, set a smoothing window with a size of smooth_size, and set the starting position to the smoothing window Half of the activity label sequence within the window, we count and sort the activities of each category, and select the activity category with the most surrounding distribution as the current activity label according to the sorting result. After updating the activity label sequence, the performance of activity recognition and segmentation. Finally, it is enhanced by linking Patches of the same activity class to determine the start and end position of each activity in the data flow. We set the boundary set of each activity $B=\{(s_{i},e_{i})\}_{i=1}^{N}$, where $s_{i}$ represents the start subscript of the i-th activity, $e_{i}$ represents the end subscript of the i-th activity, and N is the number of activities in the activity label sequence. The activity set corresponding to each boundary is represented by $\hat{a}=\{a_{i}\}_{i=1}^{N}$, where $a_{i}$ is the activity corresponding to the i-th boundary. The Process of Algorithm is described in the Algorithm 1.
III-E Activity forecast
@@ IV-A Experimental setup
Our method was implemented using PyTorch and trained on an RTX 3090 GPU. The batch size was set to 64, and we utilized the Adam optimizer with a learning rate of 0.0001.
@@ IV-B Datasets
WISDM Dataset[27]: The dataset was collected from 29 participants using mobile phones equipped with triaxial accelerometers placed in their pants pockets. The data was sampled at a frequency of 20 Hz. Participants engaged in six activities on a daily basis: walking, strolling, ascending stairs, descending stairs, standing still, and standing. To address missing values in the dataset, the mean of the corresponding column was used for filling them.
@@
PAMAP2 Dataset[28]: The dataset consists of data collected from nine participants who wore IMUs on their chest, hand, and ankle. These IMUs collected information on acceleration, angular velocity, and magnetic sensors. Each participant was instructed to complete 12 mandatory activities such as lying down, standing, going up and down stairs, as well as 6 optional activities like watching TV, driving, and playing ball.
@@
UNIMIB SHAR Dataset[29]: The dataset was collected by researchers at the University of Milan-Bicocca. Thirty volunteers, ranging in age from 18 to 60 years old, were equipped with Bosch BMA220 sensors on their Samsung mobile phones. The phones were placed in both the participants’ left and right pockets, resulting in a total of 11,771 activities recorded. The dataset comprises 17 categories of activities, categorized into two coarse-grained groups: 9 activities of daily living (ADL) and 8 falling behaviors.
IV-C Comparison algorithm
We compare our method with state-of-the-art algorithms for activity recognition, segmentation, and forecast.
  • MPCC-MRCSE[58]: A multidimensional parallel convolutional connected (MPCC) deep learning network based on MMS data for HAR is proposed that fully utilizes the advantages of multidimensional convolutional kernels.
  • CNN-BiLSTM[59]: Proposed multibranch CNN-BiLSTM network does automatic feature extraction from the raw sensor data with minimal data pre-processing
  • A&D[11]: A deep learning method for HAR that includes modules such as CIE, AGE, and center loss. The method achieves state-of-the-art recognition performance through data enhancement techniques.
  • MRFC-VAE-Net [56]: A multi-resolution fusion convolutional network is proposed to make up for the shortcoming that confusing activities are difficult to correctly identify, thereby improving the accuracy of identification.
  • Multi-ResAtt[60]: Multi-ResAtt integrates a recurrent neural network with attention to extract time-series features and perform activity recognition.
  • ConvAttention[33]: A lightweight multi-head model that incorporates an attention mechanism to enhance the representation capability of CNNs for human activity recognition.
  • CE-HAR[57]: Provide an alternative called Channel Equalization to reactivate these inhibited channels by performing whitening or decorrelation operation, which compels all channels to contribute more or less to feature representation
  • TripleCrossAttention[35]: This method introduces a triple cross-dimensional attention mechanism for sensor-based activity recognition tasks.
  • SK[36]: A multi-branch CNN model that utilizes selective kernel mechanisms for HAR. Additionally, it incorporates attention mechanisms for the first time in the HAR domain to select kernels between multiple branches with different receptive fields (RFs).
  • MTHARS[5]: A multi-task deep neural network that enhances activity segmentation and recognition by utilizing a multi-scale window splicing method. This approach effectively combines both tasks and improves performance.
  • BCMLF-HARS[6]: A multi-task learning framework for joint segmentation and recognition of activities using wearable sensors. It introduces a multi-level temporal convolutional network (MS-TCN) to predict activities at the sample level, overcoming the challenge of multi-class windows. The consistency loss attention mechanism effectively enhances the consistency between activity and boundary predictions.
IV-D Comparisons with activity recognition algorithm
@@
In this section, we compare our method with other competing methods that have been evaluated on three publicly available datasets for wearable-based Human Activity Recognition (HAR). To conduct a thorough performance analysis, we use non-overlapping patches of size 10 to compare the overall recognition performance.
@@
We compared the recognition performance of our proposed method with other similar algorithms as mentioned earlier. The recognition performance of various competing methods is summarized in Table I. Our method demonstrated improvements in recognition accuracy across three publicly available HAR datasets. Particularly, on the Unimib and PAMAP2 datasets, our method achieved the highest recognition performance, obtaining weighted average F1 scores of 82.04% and 98.92% respectively. Remarkably, our method surpassed the performance of the leading competing method by 3.26% in terms of weighted average F1 score on the Unimib dataset.
[…]
Methods Datasets PAMAP2 WISDM UNIMIB SHAR
Acc F1 Acc F1 Acc F1
P2LHAP 98.71% 98.92% 97.51% 97.52% 81.51% 82.04%
Other Reachers 90.70% 90.55% [11] 92.71% 92.63% [34] 76.30% 75.38% [35]
93.80% 93.38% [36] 96.04% 96.05% [59] 75.89% 74.63% [36]
94.27% 94.25% [59] 96.85% 97.20% [33] 76.48% 75.71% [5]
93.19% 92.96% [60] 97.51% 97.25% [36] 77.29% 76.83% [60]
94.50% 94.80% [5] 99.04% 99.18% [57] 78.65% 78.78% [57]
92.14% 92.18% [57] 97.11% 98.64% [56]
94.59% 94.47% [6]
97.39% 98.12% [56]
- 98.33% [58]
@@
IV-E Qualitative segmentation results
To demonstrate the segmentation performance of our proposed method, this section presents a visualization of patch-level activity predictions in Fig. 4, with each color representing a different activity category. We compared sequence fragments obtained through model segmentation with selected sequence fragments from the PAMAP2 and WISDM datasets. In practical application scenarios, excessive segmentation can adversely affect prediction accuracy. Nonetheless, as depicted in the figure, our method successfully reduces most over-segmentation errors, leading to more precise and consistent patch-level predicted segments within each active segment.
Table II displays the segmentation performance of our proposed method in comparison to competing algorithms. Our method excels in both segmentation and recognition tasks, consistently outperforming the comparison algorithms. Specifically, our Jaccard Index results on the Pamap2 dataset surpassed those of A&D and BCMLF-HARS by 9.82% and 3.47% respectively, Where (w/o) represents the segmentation Jaccard Index scores obtained without using the smoothing strategy.
TABLE II: RESULTS OF SEGMENTATION METRICS ON PAMAP2 DATASET.
@@ IV-E Qualitative segmentation results
Attn A&D BCMLF-HARS ours(w/o) Ours
@@
IV-F Patch prediction analysis
In this section, we assess the performance of the activity prediction module using our model to forecast the duration of activity categories in the next 8 patch blocks. Fig. 5 displays the confusion matrix for the future activity prediction in the pamap2 dataset. Fig. 5 (a),(b),(c) present the prediction confusion matrices of Conv2LSTM[38], Seq2Seq-LSTM[39], and Seq2Seq-LSTM-PE-MA[40] predictors for five activity categories in the dataset, while Fig. (d) illustrates the predictions of our proposed model. The confusion matrix results for all campaigns indicate an average accuracy of 96.59%. Additionally, we compare the predicted outputs of future sensor signals with the ground truth signals by evaluating their mean squared error (MSE). To enable the model to predict future sensor signals, we exclude the last classification header and utilize the first half of the input sensor sequence for model input, reserving the second half to match the size of the predicted output data. Table III presents the performance comparison of our model with five predictors for sensor sequence signal prediction.
@@ IV-F Patch prediction analysis
TABLE III: Evaluation results of the forecasters.
@@ IV-G Ablation experiments
Impact of Patch Size: various sizes and quantities of patches can yield different outcomes. Hence, we employ a scaling approach with multiple patches to analyze their performance. We provide the comprehensive results in Table IV. We observed a decrease in accuracy when the Patch size is reduced significantly. This decline can be attributed to the issue of excessive segmentation. However, when the Patch size is set to 10, it consistently exhibits exceptional recognition performance across all three datasets.
The impact of Patching modules: we examined the effects on the results by not patching the sensor data stream, but instead using the Transformer backbone directly. The F1 score obtained on the benchmark dataset is presented in Table V. It was observed that when the data was not patched during processing, the accuracy of the experimental results deteriorated. Hence, applying data patching enhanced the performance of activity recognition.
TABLE IV: $F_{1}$ values with different patch sizes on benchmark datasets.
Datasets Patch Size
P=1 P=2 P=5 P=10 P=20
   WISDM
0.8544 0.9056 0.8819 0.9751 0.9568
   PAMAP2
0.9151 0.9406 0.9610 0.9892 0.9667
   UNIMIB
0.6719 69.50 75.66 0.8204 74.76
[…]
@@
only transformer backbone 0.7090 0.8783 0.7533
P2LHAP $\bf{0.9752}$ $\bf{0.9892}$ $\bf{0.8204}$
V CONCLUSION AND FUTURE WORK

### v3

@@ III-B2 Encoder
We use the native Transformer encoder [14] to model the features of patch-level sensor sequences, similar to the method used in [11]. The Encoder module utilizes trainable linear projections to map continuous Patch blocks to a D-dimensional Transformer latent space $\mathbf{W}_{p}\in\mathbb{R}^{D\times P}$. It incorporates a learnable additive positional encoding $\mathbf{W}_{pos}\in\mathbb{R}^{D\times N}$ to capture the temporal order of the patches, represented by $\mathbf{x}_{d}^{(i)}=\mathbf{W}_{p}\mathbf{x}_{p}^{(i)}+\mathbf{W}_{pos}$, where $\mathbf{x}_{d}^{(i)}\in\mathbb{R}^{D\times N}$ represents the input fed into the Transformer encoder. For each head $h=1,\cdots,H$ in the multi-head attention, the Patch blocks are transformed into query matrices $\mathbf{Q}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{Q}$, key matrices $\mathbf{K}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{K}$, and value matrices $\mathbf{V}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{V}$, where $\mathbf{W}_{h}^{Q},\mathbf{W}_{h}^{K}\in\mathbb{R}^{D\times d_{k}},\mathbf{W}_{h}^{V}\in\mathbb{R}^{D\times D}$. Subsequently, the attention outputs are normalized to obtain the final attention output $\mathbf{Q}_{h}^{(i)}\in\mathbb{R}^{D\times N}:$
@@ III-B3 Decoder
$\mathbf{z}^{{}^{\prime}(i)}=softmax(\mathbf{Q}^{T}\mathbf{K}/\sigma)\mathbf{V}=f_{\theta}(\mathbf{X}^{{}^{\prime}},\mathbf{Z}^{(i)})$ (3)
Here $\sigma$ is the temperature parameter, and $f_{\theta}(\cdot)$ represents the attention mechanism network.
@@ III-C Activity classification
Among them, $\hat{y}$ and $y$ represent the patch-level predicted label and the real label, respectively. Given the imbalance of human activity classes under unconstrained conditions, we employ the Softmax CrossEntropy loss function $l_{p}(\cdot)$ and take into account the number of effective samples for classification. The calculation formula is as follows:
@@
Where $L$ signifies the length of input sequence, and C represents the activity class number of dataset. In sensor-based HAR tasks, we leverage this loss function and take into account the number of effective samples to amplify the weight assigned to minority classes. This approach aims to yield fairer and more robust results for each individual class.
@@ III-D Activity segmentation and smoothing strategy
Given that the classification loss treats each patch independently, it may result in unwarranted over-segmentation errors. To achieve smoother transitions between patches and minimize over-segmentation errors, we employ the truncated mean square error as our chosen smoothing loss function. The formulation of the smoothing loss is as follows:
@@
For the activity label sequence $\hat{y}$ output by the model, in order to reduce the impact of over-segmentation on subsequent activity merging, we set each activity label as the center, set a smoothing window with a size of smooth_size, and set the starting position to the Half of the activity label sequence within the window. Within the window, we count and sort the activities of each category, and select the activity category with the most surrounding distribution as the current activity label according to the sorting result. After updating the activity label sequence, the performance of activity recognition and segmentation. Finally, it is enhanced by linking Patches of the same activity class to determine the start and end position of each activity in the data flow. We set the boundary set of each activity $B=\{(s_{i},e_{i})\}_{i=1}^{N}$, where $s_{i}$ represents the start subscript of the i-th activity, $e_{i}$ represents the end subscript of the i-th activity, and N is the number of activities in the activity label sequence. The activity set corresponding to each boundary is represented by $\hat{a}=\{a_{i}\}_{i=1}^{N}$, where $a_{i}$ is the activity corresponding to the i-th boundary. The Process of Algorithm is described in the Algorithm 1.
III-E Activity forecasting
@@ IV-A Experimental setup
Our method was implemented using PyTorch and trained on an RTX 3090 GPU. The batch size was set to 64, and we utilized the Adam optimizer with a learning rate of 0.0001. The fully connected dropout rate is set at 20% (0.2), head dropout is fixed at 5% (0.05), the default patch length is 10, with a step size of 10. The model employs 8 heads by default, consists of 6 encoder-decoder layers, and undergoes training for 100 epochs.
@@
The choice of several parameters values for the models is shown in the table I:
TABLE I: The default choice of several parameters values for the models
      Model Parameters       Parameters Values       Parameter Introduction
      seq_len       200/151/512        input sensor sequence length 200(WISDM)/151(Unimib SHAR)/512(PAMAP2)
      pred_len       8       prediction label sequence length
      fc_dropout       0.2       fully connected dropout
      attn_dropout       0.05       attention dropout
      head_dropout       0.00       transformer head dropout
      patch_len       10       patch length
      stride       10       stride
      revin       1       RevIN True: 1 False: 0
      enc_in       3/3/9        encoder input size(the number of sensor channels) 3(WISDM)/3(Unimib SHAR)/9(PAMAP2)
      dec_in       3/3/9        decoder input size(the number of sensor channels) 3(WISDM)/3(Unimib SHAR)/9(PAMAP2)
      n_heads       8       number of transformer heads
      e_layers       6       number of encoder layers
      d_layers       6       number of decoder layers
      d_ff       256       dimension of FCN
      class_num       6/17/12       number of activities
      d_model       256       dimension of transformer model
      s_t       2       smoothing threshold
      d_k       d_model//n_heads       dimension of the query/key vectors per attention head
      d_v       d_model//n_heads       dimension of the value vectors per attention head
@@ IV-B Datasets
WISDM Dataset[30]: The data were obtained by 29 participants using phones with triaxial acceleration sensors placed in their trouser pockets with a sampling frequency of 20 Hz. Walking, strolling, walking up stairs, walking down stairs, standing motionless, and standing up were among the six daily activities undertaken by each participant. Fill missing values in a dataset using linear interpolation.
@@
PAMAP2 Dataset[31]: This dataset comprises data collected from nine participants wearing IMUs on their chest, hands, and ankles with a sampling frequency of 100 Hz. These IMUs gather information on acceleration, angular velocity, and magnetic sensors. Each participant completed 12 mandatory activities, including lying down, standing, and going up and down stairs, as well as 6 optional activities, such as watching TV, driving a car, and playing ball.
@@
UNIMIB SHAR Dataset[32]: The dataset was collected from the University of Milano-Bicocca. Thirty volunteers, aged 18 to 60, had their Samsung phones equipped with Bosch BMA 220 sensors with a sampling frequency of 50 Hz. The study recorded 11,771 activities by placing phones in the participants’ left and right pockets. The dataset comprises 17 categories of activities, divided into two groups: 9 activities of daily living (ADL) and 8 falling behaviors. Each activity was repeated 3 to 6 times.
IV-C Comparisons with activity recognition algorithm
@@
In this section, we compare our method with other competing methods that have been evaluated on three publicly available datasets for wearable-based Human Activity Recognition. To conduct a thorough performance analysis, we use non-overlapping patches of size 10 to compare the overall recognition performance.
@@
We compared the recognition performance of our proposed method with other similar algorithms as mentioned earlier. The results of various competing methods are summarized in Table II. Our method demonstrated improvements in recognition accuracy across three publicly available HAR datasets. Specifically, on the Unimib and PAMAP2 datasets, our method achieved the highest recognition performance, obtaining weighted average F1 scores of 82.04% and 98.92% respectively. Remarkably, our method surpassed the performance of the leading competing method by 3.26% in terms of weighted average F1 score on the Unimib dataset.
TABLE II: The comparison of the accuracy and F1 performance of our work with other works on three datasets.
PAMAP2 WISDM UNIMIB SHAR
methods Acc F1 methods Acc F1 methods Acc F1
ours 98.71% 98.92% ours 97.51% 97.52% ours 81.51% 82.04%
Y Li et al. (2023)[33] 98.04% - K. Xia et al. (2020)[34] 92.71% 92.63% Y. Tang et al. (2022) [35] 76.30% 75.38%
W. Gao et al. (2021)[36] 93.80% 93.38% S Kobayashi et al. (2023)[37] 90.62% - W. Gao et al. (2021) [36] 75.89% 74.63%
SK. Challa et al. (2022)[23] 94.27% 94.25% Z. N. Khan et al. (2021)[38] 96.85% 97.20% F. Duan et al. (2023)[6] 76.48% 75.71%
Al-Qaness et al. (2022)[39] 93.19% 92.96% W. Gao et al. (2021)[36] 97.51% 97.25% Al-Qaness et al. (2022)[39] 77.29% 76.83%
F. Duan et al. (2023)[6] 94.50% 94.80% W. Huang et al. (2023)[40] 99.04% 99.18% W. Huang et al. (2023)[40] 78.65% 78.78%
W. Huang et al. (2023) [40] 92.14% 92.18% J. Li et al. (2023)[41] 97.11% 98.64% Y. Tang et al. (2022) [42] 79.02% 79.19%
S. Xia et al. (2022)[7] 94.59% 94.47% SK. Challa et al. (2022)[23] 96.04% 96.05%
J. Li et al. (2023) [41] 97.39% 98.12% MK Yi et al.(2023)[17] 96.07% 96.14%
Y. Wang et al. (2023)[43] - 98.33%
@@
IV-D Qualitative segmentation results
To demonstrate our proposed method’s segmentation performance, this section presents a visualization of patch-level activity predictions in Fig. 4, with each color representing a different activity category. We compare sequence fragments obtained through model segmentation with selected sequence fragments from the PAMAP2 and WISDM datasets. In practical application scenarios, excessive segmentation can adversely affect prediction accuracy. Nonetheless, as depicted in the figure, our method successfully reduces most over-segmentation errors, leading to more precise and consistent patch-level predicted segments within each active segment.
Table III displays the segmentation performance of our proposed method in comparison with competing algorithms. Our method excels in both segmentation and recognition tasks, consistently outperforming the comparison algorithms. Notably, our Jaccard Index scores on the Pamap2 dataset surpassed those of A&D and JSR by 9.82% and 3.47% respectively, where (w/o) represents the segmentation Jaccard Index scores obtained without using the smoothing strategy. This justifies our employment of a smoothing algorithm, which effectively mitigates the boundary irregularities resulting from excessive segmentation. The algorithm serves to enhance clarity by refining the boundaries and ensures clearer and more distinct segmentation for each individual activity.
Sensor-recorded human activity data forms a continuous data stream, encompassing diverse actions like standing, sitting, walking, and transitions such as stand-to-sit and sit-to-lying. These activities are classified into two categories: basic and transition, based on their duration. Basic activities, characterized by longer periods, can be either dynamic (e.g., walking) or static (e.g., sitting), whereas transitional activities involve brief actions (seconds), like posture shifts (e.g., sit-to-stand). Despite the prevalence of research on basic activities, transition events are frequently overlooked due to their lower frequency and shorter duration compared to basic activities. When considering the short-lived activities that intervene between consecutive basic activities, two distinct scenarios may arise:
  • When two consecutive basic activities differ, the activity segments in between are classified as transition activities.
  • If adjacent basic activities are identical, the segments in between are designated as the disturbance process of the activity. This occurs due to external influences or human factors disrupting the sensor data, causing irregular fluctuations over time. Consequently, the activities preceding and following the disturbance are considered of the same type.
Given the brief nature of transition activities, our framework efficiently divides the original sensor data into multiple patches, ensuring that the information of these short events can be effectively captured and learned. As demonstrated in Fig. 4(c), the short duration of the Jogging activity does not hinder our framework’s ability to accurately segment it, showcasing its effectiveness.
TABLE III: RESULTS OF SEGMENTATION METRICS ON PAMAP2 DATASET
@@ IV-D Qualitative segmentation results
Attn A&D JSR ours(w/o) Ours
@@
IV-E Patch forecast analysis
In this section, we assess the performance of the activity forecast module using our model to forecast the duration of activity categories in the next 8 patch blocks. Fig. 5 displays the confusion matrix for the future activity forecast in the pamap2 dataset. Fig. 5 (a),(b),(c) present the prediction confusion matrices of Conv2LSTM[44], Seq2Seq-LSTM[45], and Seq2Seq-LSTM-PE-MA[46] predictors for five activity categories in the dataset, while Fig. (d) illustrates the predictions of our proposed model. The confusion matrix results for all campaigns indicate an average accuracy of 96.59%. Additionally, we compare the predicted outputs of future sensor signals with the ground truth signals by evaluating their mean squared error (MSE). To enable the model to predict future sensor signals, we exclude the last classification header and utilize the first half of the input sensor sequence for model input, reserving the second half to match the size of the predicted output data. Table IV presents the performance comparison of our model with five forecasters for sensor sequence signal prediction.
@@ IV-E Patch forecast analysis
TABLE IV: Evaluation results of the forecast on PAMAP2 dataset
@@ IV-F Ablation experiments
Impact of Patch Size: various sizes and quantities of patches can yield different outcomes. Hence, we employ a scaling approach with multiple patches to analyze their performance. We provide the comprehensive results in Table V. We observed a decrease in accuracy when the Patch size is reduced significantly. This decline can be attributed to the issue of excessive segmentation. However, when the Patch size is set to 10, it consistently exhibits exceptional recognition performance across all three datasets.
The impact of Patching modules: we examined the effects on the results by not patching the sensor data stream, but instead using the Transformer backbone directly. The F1 score obtained on the benchmark dataset is presented in Table VI. It was observed that when the data was not patched during processing, the accuracy of the experimental results deteriorated. Hence, applying data patching enhanced the performance of activity recognition.
The impact of the number of multiple heads and the number of encoder-decoder layers:this study investigates the influence of model components on classification outcomes by varying the number of multi-heads and encoder-decoder layers. Table VII presents the activity recognition accuracy of our proposed framework on the PAMAP2 dataset, showcasing the performance changes with different configurations. The results indicate that an increase in both multi-heads count and model layer depth facilitates better learning of activity sequence representations, leading to enhanced recognition accuracy. The optimal configuration is achieved when the model has 6 layers and 8 multi-heads, achieving the highest accuracy.
TABLE V: $F_{1}$ values with different patch sizes on benchmark datasets
Datasets Patch Size P=1 P=2 P=5 P=10 P=20
   WISDM 0.8544 0.9056 0.8819 0.9751 0.9568
   PAMAP2 0.9151 0.9406 0.9610 0.9892 0.9667
   UNIMIB 0.6719 0.6950 0.7566 0.8204 0.7876
TABLE VI: Activity classification F1 value without Patching module on benchmark datasets
@@
only transformer backbone 0.9090 0.8783 0.7533
P2LHAP 0.9752 0.9892 0.8204
TABLE VII: Activity recognition accuracy with different numbers of multi-heads and different numbers of encoder-decoder layers on the PAMAP2 dataset
Layer Mutil-heads 1 2 4 8 16
     1 0.8964 0.9259 0.9372 0.9447 0.9485
     2 0.9406 0.9556 0.9560 0.9654 0.9581
     3 0.9693 0.9702 0.9633 0.9750 0.9719
     6 0.9658 0.9741 0.9793 0.9871 0.9788
     8 0.9710 0.9793 0.9790 0.9802 0.9812
IV-G Further Analysis
To thoroughly assess the practicality of our proposed methodology, we will examine its computational complexity and parameter count, aiming to determine its feasibility for real-time deployment on mobile or embedded devices. The computational analysis was conducted on a Linux platform equipped with an RTX 3090 GPU and an Intel Core i9-13900K. The experiments employed a batch size of 32 for both training and testing, with 3 encoder and decoder layers, and 4 multi-heads. Utilizing the PAMAP2 dataset, we set the input sequence length to 512, patch size to 24, and step size to 12.
The computational complexity comparison, as presented in the table VIII, reveals that our method exhibits superior parameter efficiency and lower inference time compared to the benchmarked techniques.
TABLE VIII: Complexity comparison of several competing methods on the PAMAP2 dataset
Methods Param FLOPs Inference time(s/epoch)
Attn.[47] 1.983M 20.058M 8.22
A&D [48] 1.503M 25.612M 11.88
JSR [7] 2.799M 12.105M 0.78
Ours 1.449M 20.014M 0.08
V Discussion and Future Work
This section delves into the merits and drawbacks of our proposed framework, with a focus on addressing the challenges posed by the inherent uncertainty in natural human activity durations and the complexities of IMU signal analysis for activity segmentation, recognition, and forecast. To tackle these issues, we introduce a novel multi-task deep learning framework that employs the transformer’s Patch Sequence to Label Sequence architecture. The transformer network captures inter-patch relationships, enabling it to predict future activity sequences based on past representations. However, our framework’s reliance on a closed space assumption in traditional machine learning restricts its ability to handle unknown activities, as the softmax classification layer misclassifies them as known categories. Future work will concentrate on enhancing the architecture to accommodate activity recognition in novel environments, by addressing the issue of unknown activity recognition. Below we will discuss the points planned for improvement in subsequent research:
1) Given that our model employs a Softmax function for classification, its output space is confined, precluding the ability to reject unknown classes, as the predicted probabilities are normalized among all known training classes. Consequently, the model struggles to accurately identify novel activities. To address this limitation, we intend to adopt Deep Open Classification (DOC) [49] in our future work. Unlike softmax, DOC employs a "1-vs-Rest" layer with N Sigmoid functions for N known classes, which serves as a basis for representing both known and unknown classes. By transferring the learned similarity knowledge from known classes to potential unknowns, we can leverage hierarchical clustering algorithms to cluster rejected instances based on the transferred similarity, thus revealing potential activity classes within the rejected samples.
2) Our model currently adopts a supervised training approach. Future iterations will incorporate self-supervision for retraining, as it eliminates the need for costly labeled datasets in real-world activity recognition scenarios. Self-supervised learning generates training targets autonomously and exploits unlabeled data to learn representations, addressing the challenge of limited labeled data or high annotation costs. This adaptation enhances the model’s applicability to practical tasks. Moreover, by facilitating the learning of more universal and general feature representations for unseen activities, self-supervision contributes to improving the model’s generalization across diverse activity recognition tasks.
VI CONCLUSION